OpenStudy (anonymous):

THE LAST ONE: 2^8 x 4^6 x 8^8

OpenStudy (anonymous):

break every number up into prime numbers

OpenStudy (anonymous):

i worked out an answer but i'm just no sure if it's correct.

OpenStudy (anonymous):

4^6 = 2^12 8^8 = 2^24

OpenStudy (anonymous):

2^8 x(2^2)^6x(2^3^8) 2^8 x 2^12 x 2^24 2^(8+12+24) 2^44

OpenStudy (anonymous):

@jamesm and Heat: I had neither of those two answers, my answer was 2^42

OpenStudy (anonymous):

nearly right...must have made a tiny error

OpenStudy (anonymous):

oops, sorry my bad 4^22 I was looking at the wrong question D;

OpenStudy (anonymous):

yes, 4^22... which is 2^44

OpenStudy (anonymous):

I think I get it now (: thanks guys!

OpenStudy (anonymous):

4^22= (2^2)^22 = 2^44

OpenStudy (anonymous):

I think this question has been asked already.